OpenStudy (anonymous):

solve the following sytem of equation 4x+7y=-14 5x+3y=17

OpenStudy (anonymous):

you can solve this one i jsut explained how to do it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

math is all about doing the work

OpenStudy (agreene):

(7,-6)

OpenStudy (lilg132):

oh i see what i did wrong

OpenStudy (anonymous):

you have to multiply one by a negitive number to get it to cancel

OpenStudy (lilg132):

-20x - 35y = 70 20x + 12y = 68 -23y = 138 -y = 6 y = -6

OpenStudy (anonymous):

i cant get it :(

OpenStudy (lilg132):

ok let me explain :p 4x + 7y = -14 5x + 3y = 17 now we need to get rid of one variable and to do that we have to make them the same number the easiest one to get rid of is 'x'. to do this we need to make one of them -x and one +x so to get the same number for both equations we should multiply the first equation by - 5 and multiply the second equation by 4 so you will get -5(4x + 7y = -14) which will give you -20x - 35y = 70 now the second equation we dont need to change to a minus so we multiply by a positive so you will get 4(5x + 3y = 17) which will give you 20x + 12y = 68 now we need to cancel out the variable x so -20x - 35y = 70 20x +12y = 68 your left with -35y = 70 12y = 68 you add them all up which gives you -23y = 138 now to solve y you need to divide both sides by 23 so you will get -23y / 23 = 138 / 23 you will get -y = 6 now we dont want y as a minus so we just switch it around which gives you y = -6 now we know what y is so we use that to work out x. lets use it with the second equation 5x + 3y = 17 as we know what 'y' is it will look like this 5x + 3(-6) = 17 which gives you 5x - 18 = 17 now we need to rearrange it so that x is on one side and numbers on the other side remember bringing a number to the other side changes its value e.g. if it is a minus it will change to positive. so you get this 5x = 17 + 18 5x = 35 now we need to divide both sides by 5 to get x on its own 5x / 5 = 35 / 5 x = 7 :D dont forget to give me a medal